# The parametric complexity of operator learning

### v3

@@ 1. Introduction
The paper is devoted to a study of the computational complexity of the approximation of maps between Banach spaces by means of neural operators. The paper has two main focii: establishing a complexity barrier for general classes of $C^{r}-$ or Lipschitz regular maps; and then showing that this barrier can be beaten for Hamilton-Jacobi (HJ) equations. In Subsection 1.1 we give a detailed literature review; we set in context the definition of “the curse of parametric complexity” that we introduce and use in this paper; and we highlight our main contributions. Then, in Subsection 1.2, we overview the organization of the remainder of the paper.
@@ 1.1. Context and Literature Review
To overcome the general curse of parametric complexity implied by Theorem 2.11 (and Theorem 2.27), efficient operator learning frameworks therefore have to leverage additional structure present in the operators of interest, going beyond $C^{r}$- or Lipschitz-regularity. Previous work on overcoming this curse for operator learning has mostly focused on operator holomorphy [23, 50, 34] and the emulation of numerical methods [30, 34, 32, 40] as two basic mechanisms for overcoming the curse of parametric complexity for specific operators of interest. A notable exception are the complexity estimates for DeepONets in [15] which are based on explicit representation of the solution; most prominently, this is achieved via the Cole-Hopf transformation for the viscous Burgers equation.
@@
The HJ equations present an application area that has the potential to be significantly impacted by the use of ideas from neural networks, especially regarding the solution of problems for functions defined over subsets of high dimensional ($d$) Euclidean space [7, 8, 12, 11]; in particular beating the curse of dimensionality with respect to this dimension $d$ has been of focus. However, this body of work has not studied operator learning, as it concerns settings in which only fixed instances of the PDE are solved. The purpose of the second part of the present paper is to study the design and analysis of neural operators to approximate the solution operator for HJ equations; this operator maps the initial condition (a function) to the solution at a later time (another function).
@@ 2.1.2. Two Simple Facts from ReLU Neural Network Calculus
The following facts will be used without further comment (see e.g. [44, Section 2.2.3] for a discussion of more general results): If $\Psi:\mathbb{R}^{D_{\mathcal{X}}}\to\mathbb{R}^{D_{\mathcal{Y}}}$ is a ReLU neural network, and $A\in\mathbb{R}^{{D_{\mathcal{X}}}\times d}$ is a matrix, then there exists a ReLU neural network $\widetilde{\Psi}:\mathbb{R}^{d}\to\mathbb{R}^{D_{\mathcal{Y}}}$, such that
@@ 2.1.3. Approximation Theory and CoD for ReLU Networks
One overall finding of research into the approximation power of ReLU neural networks is that, for function approximation in spaces characterized by smoothness, neural networks cannot entirely overcome the curse of dimensionality [54, 55, 29, 38]. This is illustrated by the following result, which builds on [54, Thm. 5] derived by D. Yarotsky:
@@ 2.2.1. Infinite-dimensional hypercubes
The notion of infinite-dimensional cubes introduced in Definition 2.3 is only a minor generalization of an established notion of cube embeddings, introduced by Donoho [18] in a Hilbert space setting. We refer to [10, Chap. 5] for a pedagogical exposition of such cube embeddings in the Hilbert space setting, and their relation to the Kolmogorov entropy of $K$.
@@
Our definition of an infinite-dimensional hypercube is natural in view of the following lemma, the discussion following Example (2.4), and other similar results.
@@ 2.3. Main Theorem on Curse of Parametric Complexity
We finally link our main theorem to a related result for PCA-Net [32, Thm. 3.3], there derived in a complementary Hilbert space setting for $\mathcal{X}$ and $\mathcal{Y}$; the result of [32] shows that, for PCA-Net, no fixed algebraic convergence rate can be achieved in the operator learning of general $C^{r}$-operators between Hilbert spaces; this can be viewed as a milder version of the full curse of parametric complexity identified in the present work, expressed by an exponential lower complexity bound in Theorem 2.11.
@@
2.4. Examples of Oerators of Neural Network-Type
@@
$\mathrm{size}(\Psi)\geq(2p+2)^{-1}\mathrm{cmplx}(\mathcal{S}).$
Note that the dimension $p\in\mathbb{N}$ is fixed by the underlying function space $\mathcal{Y}=\mathcal{Y}(\Omega;\mathbb{R}^{p})$. Thus, Lemma 2.18 implies a lower complexity bound $\mathrm{size}(\Psi)\gtrsim\mathrm{cmplx}(\mathcal{S})$. The detailed proof is given in Appendix A.5.1. It thus follows from Corollary 2.12 that operator learning with PCA-Net suffers from a curse of parametric complexity:
@@ Discussion.
We emphasize that our notion of complexity only relates to the size of the neural network at the core of these architectures; by design, it does not take into account other factors, such as the additional complexity associated with the practical evaluation of inner products in the PCA-Net architecture, evaluation of linear encoding functionals of DeepONets, or the numerical representation of an (optimal) output PCA basis for PCA-Net or neural network basis for DeepONet. The important point is that the aforementioned factors can only increase the overall complexity of any concrete implementation; correspondingly, our proposed notion of $\mathrm{cmplx}(\mathcal{S})$, which neglects some of these additional contributions, is can be used to derive rigorous lower bounds on the overall complexity of any implementation.
@@
(2.20) $\mathrm{cmplx}(\mathcal{S}_{\epsilon}^{N_{\epsilon}})\lesssim N^{2d}\mathrm{size}(\mathcal{S}_{\epsilon}^{N_{\epsilon}}).$
As a consequence of the exponential lower bound, $\mathrm{cmplx}(\mathcal{S}_{\epsilon}^{N_{\epsilon}})\geq\exp(c\epsilon^{-1/(\alpha+1+\delta)r})$, inequality (2.20) implies that either $N_{\epsilon}$ or $\mathrm{size}(\mathcal{S}_{\epsilon}^{N_{\epsilon}})$ have to be exponentially large in $\epsilon^{-1}$, proving the claim.
We would like to point out that the additional factor depending on $N$ is natural in view of the fact that even a linear discretized FNO layer $v\mapsto Wv$, with matrix $W\in\mathbb{R}^{d_{v}\times d_{v}}$, actually corresponds to a mapping $\mathbb{R}^{N^{d}\times d_{v}}\to\mathbb{R}^{N^{d}\times d_{v}}$, $(v(x_{j_{1},\dots,j_{d}}))\mapsto(Wv(x)_{j_{1},\dots,j_{d}})$; i.e. the matrix multiplication should be viewed as being carried out in parallel at all $N^{d}$ grid points. Representing this mapping by an ordinary matrix multiplication $\mathbb{R}^{N^{d}\times d_{v}}\to\mathbb{R}^{N^{d}\times d_{v}}$ requires $N^{d}\|W\|_{0}$ degrees of freedom, and thus depends on $N$ in addition to the number of FNO parameters $\|W\|_{0}$.
For the detailed proof, we refer to Appendix A.7. ∎
@@ 3. The Hamilton-Jacobi Equation
Here $\Omega\subset\mathbb{R}^{d}$ is a $d$-dimensional domain. To simplify our analysis we consider only the case of a domain $\Omega=[0,2\pi]^{d}$ with periodic boundary conditions (so that we may identify $\Omega$ with $\mathbb{T}^{d}$.) We denote by $C^{r}_{\mathrm{per}}(\Omega)$ the space of $r$-times continuously differentiable real-valued functions with $2\pi$-periodic derivatives, with norm
@@
In equation (HJ), $u:\Omega\times[0,T]\to\mathbb{R}$, $(q,t)\mapsto u(q,t)$ is a function, depending on the spatial variable $q\in\Omega$ and on time $t\geq 0$. We will restrict attention to problems for which a classical solution $u\in C^{r}_{\mathrm{per}}(\Omega\times[0,T])$, $r\geq 2$, exists. In this setting the initial value problem (HJ) can be solved by the method of characteristics and a unique solution may be proved to exist for some $T=T(u_{0})>0.$ We may then define solution operator $\mathcal{S}_{t}^{\dagger}$ with property $u(\cdot,t)=\mathcal{S}_{t}^{\dagger}(u_{0});$ the local time of existence in time will, in general, depend on the input $u_{0}$ to $\mathcal{S}_{t}.$ The next two subsections describe, respectively, the method of characteristics and the existence of the solution operator on a time-interval $t\in[0,T]$, for all initial data from compact $\mathcal{F}$ in $C^{r}_{\mathrm{per}}(\Omega),$ for $r\geq 2.$ Thus we define $\{\mathcal{S}_{t}^{\dagger}:\mathcal{F}\subset C^{r}_{\mathrm{per}}(\Omega)\to C^{r}_{\mathrm{per}}(\Omega)\}$ for all $t\in[0,T]$, $T=T(\mathcal{F})$ sufficiently small.
We recall that throughout this paper, use of superscript $\dagger$ denotes an object defined through construction of an exact solution of (HJ), or objects used in the construction of the solution; identical notation without a superscript $\dagger$ denotes an approximation of that object.
@@ 3.1. Method of Characteristics for (HJ)
We briefly summarize this methodology; for more details see [19, Section 3.3]. Consider the following Hamiltonian system for $t\mapsto(q(t),p(t))\in\Omega\times\mathbb{R}^{d}$ defined by
@@ 3.2. Short-time Existence of $C^{r}$-solutions
We close this section on the short-time existence with the following remark.
Remark 3.5.
Classical solutions of the Hamilton-Jacobi equations can develop singularities in finite time due to the potential crossing of the spatial characteristics $q(t)$ emanating from different points; Indeed, if two spatial characteristics emanating from $q_{0}$ and $q_{0}^{\prime}$ cross in finite time, then (3.2) does not lead to a well-defined function $u(q,t)$, and the method of characteristics breaks down. Thus, our existence theorem is generally restricted to a finite time interval $[0,T^{\ast}]$. In important special cases, the crossing of characteristics is ruled out, and classical solutions exist for all time, i.e. with $T^{\ast}=\infty$. One such example is the advection equation with velocity field $v(q)$, and corresponding Hamiltonian $H(q,\nabla_{q}u)=v(q)\cdot\nabla_{q}u$, for which the complexity of operator approximation is studied computationally in [13].
@@ 4.2. Step c) Moving Least Squares
In this section, we define the interpolation mapping :$\mathcal{R}:\{q^{j}_{t},z^{j}_{t}\}_{j=1}^{N}\mapsto u({\,\cdot\,},t)$, employing reconstruction by moving least squares [53]. In general, given a function $f^{\dagger}:\Omega\to\mathbb{R}$, here assumed to be defined on the domain $\Omega=[0,2\pi]^{d}$, and given a set of (scattered) data $\{\mathfrak{q}^{j},z^{j}\}_{j=1}^{N}$ where $z^{j}=f^{\dagger}(\mathfrak{q}^{j})$ for $\mathfrak{Q}=\{\mathfrak{q}^{1},\dots,\mathfrak{q}^{N}\}\subset\Omega$, the method of moving least squares produces an approximation $f_{z,\mathfrak{Q}}\approx f^{\dagger}$, which is given by the following minimization [53, Def. 4.1]:
(4.2) $f_{z,\mathfrak{Q}}(q)=\mathrm{min}{\left\{\sum_{j=1}^{N}\left[z^{j}-P(\mathfrak{q}^{j})\right]^{2}\phi_{\delta}(q-\mathfrak{q}^{j})\,\middle|\,P\in\pi_{m}\right\}}.$
Here, $\pi_{m}$ denotes the space of polynomials of degree $m$, and $\phi_{\delta}(q)=\phi(q/\delta)$ is a compactly supported, non-negative weight function. We will assume $\phi({\,\cdot\,})$ to be smooth, supported in the unit ball $B_{1}(0)$, and positive on the ball $B_{1/2}(0)$ with radius $1/2$.
@@
Using this approximate interpolation data, let $f_{z,Q}$ be obtained by moving least squares (4.2). Then there exist constants $h_{0},\gamma,C>0$, depending only on $d$, $r$ and $\kappa\geq 1$, such that, for $h_{Q,\Omega}\leq h_{0}$ and moving least squares scale parameter $\delta:=\gamma h_{Q,\Omega}$, we have
@@
The constants $C$ and $\gamma$ in the previous proposition can be computed explicitly [53, see Thm. 3.14 and Cor. 4.8].
@@
Given the interpolation by moving least squares and the pruning algorithm, we finally define the reconstruction map :$\mathcal{R}:\{(q_{t}^{j},z_{t}^{j})\}_{j=1}^{N}\mapsto{f}_{z,Q}$ as follows:
@@
    Define $f_{z,Q}$ as the moving least squares interpolant (4.2) based on the (pruned) interpolation data $\{(q^{j_{k}}_{t},z^{j_{k}}_{t})\}_{k=1}^{m}$ with $\delta=\gamma h_{Q^{\prime},\Omega}$, and $\gamma$ the constant in Proposition 4.2.

### v4

@@ 1. Introduction
This paper is devoted to a study of the computational complexity of the approximation of maps between Banach spaces by means of neural operators. The paper has two main focii: establishing a complexity barrier for general classes of $C^{r}-$ or Lipschitz regular maps; and then showing that this barrier can be beaten for Hamilton-Jacobi (HJ) equations. In Subsection 1.1 we give a detailed literature review; we set in context the definition of “the curse of parametric complexity” that we introduce and use in this paper; and we highlight our main contributions. Then, in Subsection 1.2, we overview the organization of the remainder of the paper.
@@ 1.1. Context and Literature Review
To overcome the general curse of parametric complexity implied by Theorem 2.11 (and Theorem 2.27), efficient operator learning frameworks therefore have to leverage additional structure present in the operators of interest, going beyond $C^{r}$- or Lipschitz-regularity. Previous work on overcoming this curse for operator learning has mostly focused on operator holomorphy [30, 63, 44] and the emulation of numerical methods [39, 44, 41, 51] as two basic mechanisms for overcoming the curse of parametric complexity for specific operators of interest. A notable exception are the complexity estimates for DeepONets in [22]. These estimates are based on explicit representation of the solution; most prominently, this is achieved via the Cole-Hopf transformation for the viscous Burgers equation. Another distinct mechanism to overcome the curse of parametric complexity is highlighted by the notion of Barron spaces, e.g. [5, 4, 50], which have recently been extended to the operator learning setting in [38]. In the finite-dimensional context, functions belonging to such Barron spaces can be approximated at Monte-Carlo rates which are independent of the dimension of the underlying domain. As shown in [38], similar rates can be established in an infinite-dimensional Barron class setting. The curse of parametric complexity derived in the present work indicates that spaces of $r$-times Fréchet differentiable operators cannot embed into operator Barron spaces, no matter the degree of differentiability $r$.
@@
The HJ equations present an application area that has the potential to be significantly impacted by the use of ideas from neural networks, especially regarding the solution of problems for functions defined over subsets of high dimensional ($d$) Euclidean space [12, 13, 19, 17]; in particular beating the curse of dimensionality with respect to this dimension $d$ has been of focus. We highlight work which proposes adapted neural network architectures, reflecting known structure of viscosity solutions of the HJ equation, to approximate and represent individual solutions [18, 52]. This cited research includes studies of analytical convergence rates for such architectures, as well as empirical work studying their practical performance [55, 34, 16]. However, this entire body of work has not studied operator learning, as it concerns settings in which only fixed instances of the PDE are solved. The purpose of the second part of the present paper is to study the design and analysis of neural operators to approximate the solution operator for HJ equations; this operator maps the initial condition (a function) to the solution at a later time (another function).
@@ 2.1.2. Two Simple Facts from ReLU Neural Network Calculus
The following facts will be used without further comment (see e.g. [57, Section 2.2.3] for a discussion of more general results): If $\Psi:\mathbb{R}^{D_{\mathcal{X}}}\to\mathbb{R}^{D_{\mathcal{Y}}}$ is a ReLU neural network, and $A\in\mathbb{R}^{{D_{\mathcal{X}}}\times d}$ is a matrix, then there exists a ReLU neural network $\widetilde{\Psi}:\mathbb{R}^{d}\to\mathbb{R}^{D_{\mathcal{Y}}}$, such that
@@ 2.1.3. Approximation Theory and CoD for ReLU Networks
One overall finding of research into the approximation power of ReLU neural networks is that, for function approximation in spaces characterized by smoothness, neural networks cannot entirely overcome the curse of dimensionality [67, 68, 37, 48]. This is illustrated by the following result, which builds on [67, Thm. 5] derived by D. Yarotsky:
@@ 2.2.1. Infinite-dimensional hypercubes
The notion of infinite-dimensional cubes introduced in Definition 2.3 is only a minor generalization of an established notion of cube embeddings, introduced by Donoho [25] in a Hilbert space setting. We refer to [15, Chap. 5] for a pedagogical exposition of such cube embeddings in the Hilbert space setting, and their relation to the Kolmogorov entropy of $K$.
@@
Our definition of an infinite-dimensional hypercube is natural in view of the following lemma, the discussion following Example 2.2, and other similar results.
@@ 2.3. Main Theorem on Curse of Parametric Complexity
We finally link our main theorem to a related result for PCA-Net [41, Thm. 3.3], there derived in a complementary Hilbert space setting for $\mathcal{X}$ and $\mathcal{Y}$; the result of [41] shows that, for PCA-Net, no fixed algebraic convergence rate can be achieved in the operator learning of general $C^{r}$-operators between Hilbert spaces; this can be viewed as a milder version of the full curse of parametric complexity identified in the present work, expressed by an exponential lower complexity bound in Theorem 2.11.
@@
2.4. Examples of Operators of Neural Network-Type
@@
$\mathrm{size}(\Psi)\geq(2p+2)^{-1}\mathrm{cmplx}(\mathcal{S}),$
where the dimension $p\in\mathbb{N}$ is fixed by the output-function space $\mathcal{Y}=\mathcal{Y}(\Omega;\mathbb{R}^{p})$.
Thus, Lemma 2.18 implies a lower complexity bound $\mathrm{size}(\Psi)\gtrsim\mathrm{cmplx}(\mathcal{S})$. The detailed proof is given in Appendix A.5.1. It thus follows from Corollary 2.12 that operator learning with PCA-Net suffers from a curse of parametric complexity:
@@ Discussion.
We emphasize that our notion of complexity only relates to the size of the neural network at the core of these architectures; by design, it does not take into account other factors, such as the additional complexity associated with the practical evaluation of inner products in the PCA-Net architecture, evaluation of linear encoding functionals of DeepONets, or the numerical representation of an (optimal) output PCA basis for PCA-Net or neural network basis for DeepONet. The important point is that the aforementioned factors can only increase the overall complexity of any concrete implementation; correspondingly, our proposed notion of $\mathrm{cmplx}(\mathcal{S})$, which neglects some of these additional contributions, can be used to derive rigorous lower bounds on the overall complexity of any implementation.
@@
(2.20) $\mathrm{cmplx}(\mathcal{S}_{\epsilon}^{N_{\epsilon}})\lesssim N_{\epsilon}^{2d}\mathrm{size}(\mathcal{S}_{\epsilon}^{N_{\epsilon}}).$
As a consequence of the exponential lower bound, $\mathrm{cmplx}(\mathcal{S}_{\epsilon}^{N_{\epsilon}})\geq\exp(c\epsilon^{-1/(\alpha+1+\delta)r})$, inequality (2.20) implies that either $N_{\epsilon}$ or $\mathrm{size}(\mathcal{S}_{\epsilon}^{N_{\epsilon}})$ have to be exponentially large in $\epsilon^{-1}$, proving the claim. For the detailed proof, we refer to Appendix A.7. ∎
Remark 2.28.
We would like to point out that the additional factor in (2.20), depending on $N_{\epsilon}$, is natural in view of the fact that even a linear discretized FNO layer $v\mapsto Wv$, with matrix $W\in\mathbb{R}^{d_{v}\times d_{v}}$, actually corresponds to a mapping $\mathbb{R}^{N_{\epsilon}^{d}\times d_{v}}\to\mathbb{R}^{N_{\epsilon}^{d}\times d_{v}}$, $(v(x_{j_{1},\dots,j_{d}}))\mapsto(Wv(x)_{j_{1},\dots,j_{d}})$; i.e. the matrix multiplication should be viewed as being carried out in parallel at all $N_{\epsilon}^{d}$ grid points. Representing this mapping by an ordinary matrix multiplication $\mathbb{R}^{N_{\epsilon}^{d}\times d_{v}}\to\mathbb{R}^{N_{\epsilon}^{d}\times d_{v}}$ requires $N_{\epsilon}^{d}\|W\|_{0}$ degrees of freedom, and thus depends on $N_{\epsilon}$ in addition to the number of FNO parameters $\|W\|_{0}$.
@@ 3. The Hamilton-Jacobi Equation
To simplify our analysis we consider only the case of a domain $\Omega=[0,2\pi]^{d}$ with periodic boundary conditions (so that we may identify $\Omega$ with $\mathbb{T}^{d}$.) We denote by $C^{r}_{\mathrm{per}}(\Omega)$ the space of $r$-times continuously differentiable real-valued functions with $2\pi$-periodic derivatives, with norm
@@
In equation (HJ), $u:\Omega\times[0,T]\to\mathbb{R}$, $(q,t)\mapsto u(q,t)$ is a function, depending on the spatial variable $q\in\Omega$ and on time $t\geq 0$. We will restrict attention to problems for which a classical solution $u\in C^{r}_{\mathrm{per}}(\Omega\times[0,T])$, $r\geq 2$, exists. In this setting the initial value problem (HJ) can be solved by the method of characteristics and a unique solution may be proved to exist for some $T=T(u_{0})>0.$ We may then define solution operator $\mathcal{S}_{t}^{\dagger}$ with property $u(\cdot,t)=\mathcal{S}_{t}^{\dagger}(u_{0});$ the maximal time $T$ of existence will, in general, depend on the input $u_{0}$ to $\mathcal{S}_{t}.$ The next two subsections describe, respectively, the method of characteristics and the existence of the solution operator on a time-interval $t\in[0,T]$, for all initial data from compact $\mathcal{F}$ in $C^{r}_{\mathrm{per}}(\Omega),$ for $r\geq 2.$ Thus we define $\{\mathcal{S}_{t}^{\dagger}:\mathcal{F}\subset C^{r}_{\mathrm{per}}(\Omega)\to C^{r}_{\mathrm{per}}(\Omega)\}$ for all $t\in[0,T]$, $T=T(\mathcal{F})$ sufficiently small.
We recall that throughout this paper, use of a superscript $\dagger$ denotes an object defined through construction of an exact solution of (HJ), or objects used in the construction of the solution; identical notation without a superscript $\dagger$ denotes an approximation of that object.
@@ 3.1. Method of Characteristics for (HJ)
We briefly summarize this methodology; for more details see [26, Section 3.3]. Consider the following Hamiltonian system for $t\mapsto(q(t),p(t))\in\Omega\times\mathbb{R}^{d}$ defined by
@@ 3.2. Short-time Existence of $C^{r}$-solutions
3.3. Limitations of the setting considered in this work
In this paper we study the curse of parametric complexity for operator learning, and discuss settings where it can be overcome. The primary role of our study of Hamilton-Jacobi equations is to highlight a setting where operator learning does not suffer from this curse. However our analysis is restricted to the setting of smooth solutions and non-intersecting characteristics. We close this section on Hamilton-Jacobi equations by highlighting this important limitation of the present work, and mention a possible extension.
The main limitation of this work is the assumption that classical solutions of (HJ) exist. Indeed, it is well-known, e.g. [1, 2], that classical solutions of the Hamilton-Jacobi equations can develop singularities in finite time due to the potential crossing of the spatial characteristics $q(t)$ emanating from different points; indeed, if two spatial characteristics emanating from $q_{0}$ and $q_{0}^{\prime}$ cross in finite time, then (3.2) does not lead to a well-defined function $u(q,t)$, and the method of characteristics breaks down. Thus, our existence theorem is generally restricted to a finite time interval $[0,T^{\ast}]$.
In certain special cases, the crossing of characteristics is ruled out, and classical solutions exist for all time, i.e. with $T^{\ast}=\infty$. One such example is the advection equation with velocity field $v(q)$, and corresponding Hamiltonian $H(q,\nabla_{q}u)=v(q)\cdot\nabla_{q}u$, for which the complexity of operator approximation is studied computationally in [20].
Finding more general conditions for the existence of classical solutions is a delicate question, which is discussed in [6]; under certain convexity assumptions on the Hamiltonian, it is possible to relate the existence of a classical solution of (HJ) to its time-reversibility. For example, it is shown in [6, Thm. 2.5] that if $H(q,p)$ is (i) strictly convex and superlinear in $p$, and (ii) Lipschitz in $q$, and $u(x,t)$ is a locally Lipschitz continuous (viscosity) solution of (HJ) both forward and backward in time, then $u$ must be smooth. Under certain technical assumptions on the convexity of $H$ and the initial data $u(t=0)$, it can in fact be shown [6] that a forward solution $u(x,t)$ and a backward solution $w(x,t)$ of (HJ) must be equal everywhere, provided that they coincide at the endpoints, i.e. $u(x,T)=w(x,T)$ and $u(x,0)=w(x,0)$Rightarrow $u\equiv w$ and $u,w\in C^{1}([0,T]\times\mathbb{R}^{d})$.
The existence of classical characteristics will play a fundamental role in the neural-network based approach to be developed in the next section. An interesting question for future work is whether classical characteristics could be replaced by a suitable notion of generalized characteristics, as studied in [1, 2]. The main challenge in such an extension is the non-uniqueness of generalized characteristics; it is not immediately clear whether a neural network-based approach can coherently approximate such generalized characteristics when there is a priori no flow map $(q_{t},p_{t})=\phi_{t}(q_{0},p_{0})$ in the classical sense. In contrast, if such a flow map exists, then this represents sufficient structure to beat the curse of parametric complexity, as shown in the following section.
@@ 4.2. Step c) Moving Least Squares
In this section, we define the interpolation mapping :$\mathcal{R}:\{q^{j}_{t},z^{j}_{t}\}_{j=1}^{N}\mapsto u({\,\cdot\,},t)$, employing reconstruction by moving least squares [66]. In general, given a function $f^{\dagger}:\Omega\to\mathbb{R}$, here assumed to be defined on the domain $\Omega=[0,2\pi]^{d}$, and given a set of (scattered) data $\{\mathfrak{q}^{j},z^{j}\}_{j=1}^{N}$ where $z^{j}=f^{\dagger}(\mathfrak{q}^{j})$ for $\mathfrak{Q}=\{\mathfrak{q}^{1},\dots,\mathfrak{q}^{N}\}\subset\Omega$, the method of moving least squares produces an approximation $f_{z,\mathfrak{Q}}\approx f^{\dagger}$, which is given by the following minimization [66, Def. 4.1]:
(4.2) $f_{z,\mathfrak{Q},n}(q)=\mathrm{min}{\left\{\sum_{j=1}^{N}\left[z^{j}-P(\mathfrak{q}^{j})\right]^{2}\phi_{\delta}(q-\mathfrak{q}^{j})\,\middle|\,P\in\pi_{n}\right\}}.$
Here, $\pi_{n}$ denotes the space of polynomials of degree $n$, and $\phi_{\delta}(q)=\phi(q/\delta)$ is a compactly supported, non-negative weight function. In the following, we will choose $n:=r-1$, where $r$ is the degree of smoothness of $f^{\dagger}\in C^{r}(\Omega)$. We will assume $\phi({\,\cdot\,})$ to be smooth, supported in the unit ball $B_{1}(0)$, and positive on the ball $B_{1/2}(0)$ with radius $1/2$.
@@
Using this approximate interpolation data, let $f_{z,Q}:=f_{z,Q,n}$ be obtained by moving least squares (4.2) with $n:=r-1$. Then there exist constants $h_{0},\gamma,C>0$, depending only on $d$, $r$ and $\kappa\geq 1$, such that, for $h_{Q,\Omega}\leq h_{0}$ and moving least squares scale parameter $\delta:=\gamma h_{Q,\Omega}$, we have
@@
The constants $C$ and $\gamma$ in the previous proposition can be computed explicitly [66, see Thm. 3.14 and Cor. 4.8].
@@
Given the interpolation by moving least squares and the pruning algorithm, we finally define the reconstruction map :$\mathcal{R}:\{(q_{t}^{j},z_{t}^{j})\}_{j=1}^{N}\mapsto{f}_{z,Q}$. In our application, the interpolation data is connected with a ground truth map $f^{\dagger}\in C^{r}(\Omega)$. The following algorithm assumes knowledge of the degree of smoothness $r$:
@@
    Given a degree of smoothness $r\in\mathbb{N}_{\geq 2}$, define $f_{z,Q}:=f_{z,Q,n=r-1}$ as the moving least squares interpolant (4.2), based on the (pruned) interpolation data $\{(q^{j_{k}}_{t},z^{j_{k}}_{t})\}_{k=1}^{m}$ with $\delta=\gamma h_{Q^{\prime},\Omega}$, and $\gamma$ the constant in Proposition 4.2.